# Controlling the Flow: Stability and Convergence for Stochastic Gradient Descent with Decaying Regularization

### v1

@@ C.1 General convergence result
Theorem C.3 (Almost sure convergence).
Suppose that 1.1 and 1.2 are fulfilled and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with predictable (random) step-sizes and regularization parameters that are uniformly bounded from above. Moreover, we assume that almost surely $(\lambda_{k})_{k\in\mathbb{N}}$ is decreasing to 00 and
@@
where we used 1.2 and Lemma B.1 (iii). Strong convexity of $f_{\lambda_{k}}$ yields
@@
Now, recall that Proposition C.1 gives that for all $k\in\mathbb{N}$
$\mathbb{E}[1\mathrm{l}_{\mathbb{A}_{k}}E_{k}\mid\mathcal{F}_{k-1}]$ $\leq\Big{(}1-2\lambda_{k}\alpha_{k}\Bigl{(}1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Bigr{)}+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}A\Big{)}1\mathrm{l}_{\mathbb{A}_{k}}E_{k-1}$
@@
where $E_{k}=f_{\lambda_{k+1}}(X_{k})-f_{\lambda_{k+1}}(x_{\lambda_{k+1}})$. Fix $N\in\mathbb{N}$ and for $k\geq N$ denote $\mathbb{B}_{k}(N)=\{\alpha_{i}\leq\frac{2}{L+\lambda_{i}}:i=N,\dots,k\}$. Then, for all $k>N$
$\begin{split}\mathbb{E}[1\mathrm{l}_{\mathbb{B}_{k}(N)}(\varphi_{k}+E_{k})\mid\mathcal{F}_{k-1}]&\leq\tau_{k}1\mathrm{l}_{\mathbb{B}_{k-1}(N)}(\varphi_{k-1}+E_{k-1})+\alpha_{k}\lambda_{k}(\|x_{\ast}\|_{\mathcal{X}}^{2}-\|x_{\lambda_{k}}\|_{\mathcal{X}}^{2})\\ &+\frac{\lambda_{k}-\lambda_{k+1}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}+\alpha_{k}^{2}\Bigl{(}A\frac{\lambda_{k}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}+C\Bigr{)}\Bigl{(}1+\frac{L+\lambda_{k}}{2}\Bigr{)},\end{split}$ (18)
@@
and we have used that $\varphi_{k-1}+E_{k-1}\geq 0$ and $B_{k-1}(N)\supset B_{k}(N)$. Now, we can apply Corollary B.8 for the process $(1\mathrm{l}_{\mathbb{B}_{k}(N)}(\varphi_{k}+E_{k}))_{k\geq N}$ to deduce that, on $B_{\infty}(N)=\bigcap_{k\geq N}B_{k}(N)$, one has $\varphi_{k}\to 0$ almost surely as $k\to\infty$. Since $\alpha_{k}\to 0$ almost surely one has
$\mathbb{P}\Big{(}\bigcup_{N\in\mathbb{N}}B_{\infty}(N)\Big{)}=1$
@@
Theorem C.4 ($L^{2}$-convergence).
Suppose that 1.1 and 1.2 are fulfilled and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with decreasing $(\lambda_{k})_{k\in\mathbb{N}}$. Moreover, assume that $\lambda_{k}\to 0$ and (15), or, alternatively,
@@
First, we prove the theorem assuming that (15) holds. By assumption, one has $\sum_{k\in\mathbb{N}}(1-\tau_{k})=\infty$, $\sum_{k\in\mathbb{N}}\alpha_{k}^{2}<\infty$, $\sum_{k\in\mathbb{N}}\alpha_{k}\lambda_{k}(\|x_{\ast}\|_{\mathcal{X}}^{2}-\|x_{\lambda_{k}}\|_{\mathcal{X}}^{2})<\infty$ and $\sum_{k=1}^{\infty}(\lambda_{k}-\lambda_{k+1})=\lambda_{1}<\infty$. Therefore, after taking expectations in (18), we can apply Lemma B.5 to deduce that $\mathbb{E}[\varphi_{k}]\to 0$ and $\mathbb{E}[E_{k}]\to 0$.
@@
C.2 Deterministic case: Convergence rate for reg-GD
In the following, we quantify the rate of convergence of reg-GD defined in (3). For sufficiently small step-sizes, i.e. $\alpha_{k}\leq\frac{2}{L}$ , the results are consistent with the findings for the solution of the regularized steepest descent ODE
$\dot{x}(t)+\nabla f(x(t))+\lambda(t)x(t)$
@@ C.2 Deterministic case: Convergence rate for reg-GD
derived in [7, Theorem 5] when defining the numerical time $t_{k}=\sum_{i=1}^{k}\alpha_{i}$ for $k\in\mathbb{N}$ and noting that $t_{k}\sim\frac{C_{\alpha}}{1-q}k^{1-q}$ and $\lambda_{k}\sim C_{\lambda}(\frac{1-q}{C_{\alpha}}t_{k})^{-p/(1-q)}$ for $q<1$. The proof follows the strategy of [7, Theorem 5].
Theorem C.5.
@@
Suppose that 1.1 is satisfied. Let $C_{\alpha},C_{\lambda}>0$, $p\in(0,1]$ and $q\in[0,1-p]$. Let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (3) for all $k\in\mathbb{N}$, $(\lambda_{k})_{k\in\mathbb{N}}=(C_{\lambda}k^{-p})_{k\in\mathbb{N}}$ and $(\alpha_{k})_{k\in\mathbb{N}}=(C_{\alpha}k^{-q})_{k\in\mathbb{N}}$ such that the following conditions are satisfied:
@@
(i): Proposition C.1 with $D_{k}\equiv 0$ guarantees that, for all $k\in\mathbb{N}_{0}$ with $\alpha_{k}\leq\frac{2}{L+\lambda_{k}}$, one has
@@
Set $\beta=1-q$ and for $k\in\mathbb{N}$ define [^1] In order to avoid confusion, we note that $\varphi_{k}$ is defined differently as in the proofs of Theorem C.3 and Theorem C.4 $\varphi_{k}=E_{k}k^{\beta}$. By assumption on $(\alpha_{k})_{k\in\mathbb{N}}$ one has $\alpha_{k}<\frac{2}{L}$ for all but finitely many indices $k$. Therefore there exists an $N\in\mathbb{N}$ such that for all $k>N$
@@
where in the case $q=1-p$ and $q\neq 0$ we have used that $2C_{\lambda}C_{\alpha}>1-q$ and in the case $q=0$ and $p=1$ we have used that $12C_{\lambda}C_{\alpha}(1-\mbox{$\frac{LC_{\alpha}}{2}$})>$ 1. Inserting these inequalities in (21),
@@
(ii): Follows from (i) and Lemma C.2, using that $p\leq 1-q$.
(iii): Follows from (i) and Lemma C.2.
@@
In the spirit of Section 2.3, we will derive optimal decay rates for the step-size and regularization decay for the convergence to the minimum-norm solution under the additional assumption that there exist $C_{\text{reg}},\xi>0$ with
@@
see also Section 3. Using Theorem C.5, one has
@@
Figure 12: Convergence rate for $\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}$ in the situation of Theorem C.5 under the Polyak-Łojasiewicz inequality, i.e. under (9) with $\xi=\frac{1}{4}$.
C.3 $L^{2}$-convergence rate for reg-SGD: Proof of Theorem 2.3
In the following, we formulate Theorem 2.3 in more details and provide a full prove.
Theorem C.6.
Suppose that 1.1 and 1.2 are satisfied. Let $C_{\alpha},C_{\lambda}>0$, $p\in(0,\frac{1}{2}]$ and $q\in(p,1-p]$. Let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with $(\alpha_{k})_{k\in\mathbb{N}}=(C_{\alpha}k^{-q})_{k\in\mathbb{N}}$ and $(\lambda_{k})_{k\in\mathbb{N}}=(C_{\lambda}k^{-p})_{k\in\mathbb{N}}$. If $q=1-p$ we additionally assume that $2C_{\lambda}C_{\alpha}>1-q$. Then, it holds that $\lim_{k\to\infty}\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]=0$ and
@@
We will only prove the first claim (i). Due to Theorem 2.2, one has $\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]\to 0$ if $q>p$ and $p+q<1$. The statements (ii)-(iii) follow analogously to the proof of Theorem C.5. Let $\beta=\min(1-q,q-p)$ and $\varphi_{k}=\mathbb{E}[E_{k}]k^{\beta}$. By assumption on $(\alpha_{k})_{k\in\mathbb{N}}$ one has $\alpha_{k}<\frac{2}{L}$ for all but finitely many $k\in\mathbb{N}$ so that, using Proposition C.1,
@@
Since $q>p$, $p+q\geq 1$, and $2C_{\lambda}C_{\alpha}>\min(p,1-2p)$ in the case that $q=1-p$ one can show as in (22) that there exist $\varepsilon^{\prime}>0$ and $N\in\mathbb{N}$ such that for all $k\geq N$
@@
By choice of $\beta$, one has $p+q=\max(1+p-\beta,2q-\beta)$. Therefore, we can show analogously to the proof of Theorem C.5 that
@@
C.4 Almost sure convergence rate for reg-SGD: Proof of Theorem 2.4
In the following, we formulate Theorem 2.4 in more details and provide a full prove. The proof of the almost sure convergence rates requires a sophisticated application of the Robbins-Siegmund theorem, Corollary B.7. For this, we use the variation of constants formula to separate the influence of the stochastic noise term $(D_{k})_{k\in\mathbb{N}}$ and the deterministic change in the global minimum of the regularized objective function $(x_{\lambda_{k}}-x_{\lambda_{k+1}})_{k\in\mathbb{N}}$.
Theorem C.7.
Suppose that 1.1 and 1.2 are satisfied. Let $C_{\alpha},C_{\lambda}>0$, $p\in(0,\frac{1}{2})$ and $q\in(\frac{1}{2},1-p]$. Let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with $(\alpha_{k})_{k\in\mathbb{N}}=(C_{\alpha}k^{-q})_{k\in\mathbb{N}}$ and $(\lambda_{k})_{k\in\mathbb{N}}=(C_{\lambda}k^{-p})_{k\in\mathbb{N}}$. Let $\beta\in(0,2q-1)$ and, if $q=1-p$, we assume that $2C_{\lambda}C_{\alpha}>\beta$. Then,
@@
We will only prove property (i). Properties (ii)-(iv) follow analogously to the proof of Theorem C.5. By Proposition C.1, for all $k\in\mathbb{N}$ with $\alpha_{k}\leq\frac{2}{L+\lambda_{k}}$ one has
@@
Let $\beta\in(0,2q-1)$ and $\varphi_{k}=\tilde{E}_{k}k^{\beta}$. By assumption on $(\alpha_{k})_{k\in\mathbb{N}}$ one has $\alpha_{k}<\frac{2}{L}$ for all but finitely many $k\in\mathbb{N}$ so that
$\mathbb{E}[\varphi_{k}\mid\mathcal{F}_{k-1}]$ $\leq\Big{(}1-2\lambda_{k}\alpha_{k}\Big{(}1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big{)}+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}A\Big{)}\frac{k^{\beta}}{(k-1)^{\beta}}\varphi_{k}$
@@
Since $q>p$, $p+q\geq 1$, and $2C_{\lambda}C_{\alpha}>\beta$ in the case that $q=1-p$, one can show as in (22) that there exist $\varepsilon^{\prime}>0$ and $N\in\mathbb{N}$ such that for all $k\geq N$
@@
In order to apply the Robbins-Siegmund theorem, Corollary B.7, we first prove that $(\Psi_{k})_{k\in\mathbb{N}}$ converges to 00 and, thus, is bounded. Moreover, we derive a rate of convergence for $(\Psi_{k})_{k\in\mathbb{N}}$.
Note that $(\Psi_{k})_{k\in\mathbb{N}}$ is a deterministic sequence that satisfies for all $k\in\mathbb{N}$
$\Psi_{k}=\Big{(}1-2\lambda_{k}\alpha_{k}\Big{(}1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big{)}+\frac{L+\lambda_{k}}{2}\alpha_{j}^{2}A\Big{)}\Psi_{k-1}+\frac{\lambda_{k}-\lambda_{k+1}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}.$
Therefore, analogously to the proof in the deterministic setting, see Theorem C.5, we get $\Psi_{k}\in\mathcal{O}(k^{-1+q})$. This implies that $(\Psi_{k})_{k\in\mathbb{N}}$ is bounded and, subsequently, $(\tilde{E}_{k})_{k\in\mathbb{N}}$ is almost surely bounded from below. Now, since $\beta<2q-1$ we get $\sum k^{-2q+\beta}<\infty$. Therefore, we can apply Corollary B.7 to get almost sure convergence of $(\varphi_{k})_{k\in\mathbb{N}}$ and, thus, $\tilde{E}_{k}\in\mathcal{O}(k^{-\beta})$ almost surely. Together with $\tilde{E}_{k}=E_{k}-\Psi_{k}$ and $\Psi_{k}=\mathcal{O}(k^{-1+q})$, this implies that $E_{k}\in\mathcal{O}(k^{-\min(\beta,1-q)})$ almost surely. ∎
Appendix D Properties of the Tikhonov regularization

### v2

@@ D.1 General convergence result
Theorem D.3 (Almost sure convergence).
Suppose that Section˜1.1 and Section˜1.1 are fulfilled and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with predictable (random) step-sizes and regularization parameters that are uniformly bounded from above. Moreover, we assume that almost surely $(\lambda_{k})_{k\in\mathbb{N}}$ is decreasing to 0 and
@@
where we used Section˜1.1 and Appendix˜B (iii). Strong convexity of $f_{\lambda_{k}}$ yields
@@
Now, recall that Appendix˜D gives that for all $k\in\mathbb{N}$
$\mathbb{E}[1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}E_{k}\mid\mathcal{F}_{k-1}]$ $\leq\Big(1-2\lambda_{k}\alpha_{k}\Bigl(1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Bigr)+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}A\Big)1\hskip-2.78857pt\mathrm{l}_{\mathbb{A}_{k}}E_{k-1}$
@@
where $E_{k}=f_{\lambda_{k+1}}(X_{k})-f_{\lambda_{k+1}}(x_{\lambda_{k+1}})$. Fix $N\in\mathbb{N}$ and for $k\geq N$ denote $\mathbb{B}_{k}(N)=\{\alpha_{i}\leq\frac{1}{L+\lambda_{i}}:i=N,\dots,k\}$. Then, for all $k>N$
$\begin{split}\mathbb{E}[1\hskip-2.78857pt\mathrm{l}_{\mathbb{B}_{k}(N)}(\varphi_{k}+E_{k})\mid\mathcal{F}_{k-1}]&\leq\tau_{k}1\hskip-2.78857pt\mathrm{l}_{\mathbb{B}_{k-1}(N)}(\varphi_{k-1}+E_{k-1})+\alpha_{k}\lambda_{k}(\|x_{\ast}\|_{\mathcal{X}}^{2}-\|x_{\lambda_{k}}\|_{\mathcal{X}}^{2})\\ &+\frac{\lambda_{k}-\lambda_{k+1}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}+\alpha_{k}^{2}\Bigl(A\frac{\lambda_{k}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}+C\Bigr)\Bigl(1+\frac{L+\lambda_{k}}{2}\Bigr),\end{split}$ (20)
@@
and we have used that $\varphi_{k-1}+E_{k-1}\geq 0$ and $\mathbb{B}_{k-1}(N)\supset\mathbb{B}_{k}(N)$. On the event $\mathbb{B}_{k-1}(N)$ we have
$\tau_{k}\leq 1-\underbrace{\alpha_{k}\lambda_{k}}_{=:A_{k}}+\underbrace{\Bigl(\frac{L+\lambda_{k}}{2}A+A+2L+2\lambda_{k}\Bigr)\alpha_{k}^{2}}_{=:C_{k}}\,$, (21)
where by assumption $\sum_{k\in\mathbb{N}}C_{k}<\infty$ and $\sum_{k\in\mathbb{N}}A_{k}=\infty$ almost surely. Now, we can apply Appendix˜B for the process $(1\hskip-2.78857pt\mathrm{l}_{\mathbb{B}_{k}(N)}(\varphi_{k}+E_{k}))_{k\geq N}$ to deduce that, on $\mathbb{B}_{\infty}(N)=\bigcap_{k\geq N}\mathbb{B}_{k}(N)$, one has $\varphi_{k}\to 0$ almost surely as $k\to\infty$. Since $\alpha_{k}\to 0$ almost surely one has
$\mathbb{P}\Big(\bigcup_{N\in\mathbb{N}}\mathbb{B}_{\infty}(N)\Big)=1$
@@
Theorem D.4 ($L^{2}$-convergence).
Suppose that Section˜1.1 and Section˜1.1 are fulfilled and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with deterministic step-sizes and deterministic and decreasing regularization parameters $(\lambda_{k})_{k\in\mathbb{N}}$. Moreover, assume that $\lambda_{k}\to 0$ and (17), or, alternatively,
@@
First, we prove the theorem assuming that (17) holds. By assumption, one has $\sum_{k\in\mathbb{N}}A_{k}=\infty$, $\sum_{k\in\mathbb{N}}C_{k}<\infty$, $\sum_{k\in\mathbb{N}}\alpha_{k}^{2}<\infty$, $\sum_{k\in\mathbb{N}}\alpha_{k}\lambda_{k}(\|x_{\ast}\|_{\mathcal{X}}^{2}-\|x_{\lambda_{k}}\|_{\mathcal{X}}^{2})<\infty$ and $\sum_{k=1}^{\infty}(\lambda_{k}-\lambda_{k+1})=\lambda_{1}<\infty$, where $(A_{k})$ and $(C_{k})$ are defined in (21). Therefore, after taking expectations in (20), we can apply Appendix˜B for the deterministic process $(Y_{k})_{k\in\mathbb{N}}=(\mathbb{E}[\varphi_{k}+E_{k}])_{k\in\mathbb{N}}$ to deduce that $\mathbb{E}[\varphi_{k}]\to 0$ and $\mathbb{E}[E_{k}]\to 0$.
@@
D.2 Deterministic case: Convergence rate for reg-GD
Before discussing the convergence rates for reg-GD, we want to relate our analysis to the literature in convex optimization. For this purpose we formulate our task of finding the minimum-norm solution as constrained optimization problem in form of
$\min_{x\in\mathcal{X}}\frac{1}{2}\|x\|_{\mathcal{X}}^{2}\quad\text{s.t.}\quad x\in C:=\arg\min_{y\in\mathcal{X}}f(y)\,.$
This naturally relates to the task of solving general variational inclusions of form
$0\in A(x)+N_{C}(x)$
where $A$ denotes a (maximal) monotone operator and $N_{C}(x)=\{v\in\mathcal{X}:\langle v,w-x\rangle\leq 0\quad\forall w\in C\}$ is the normal cone of a closed convex set $C$ at $x$. In our setting the operator $A(x)=\nabla_{z}\frac{1}{2}\|z\|_{\mathcal{X}}^{2}\mid_{z=x}=x$ is strongly monotone. Another important class of problems studied in this context are hierarchical optimization problems of finding points in the set
$S=\operatorname*{arg\,min}\{g(x)\mid x\in\operatorname*{arg\,min}f(x)\}$
for two convex functions $g$ and $f$. This relates to our setting by choosing $g(\cdot)=\|\cdot\|_{\mathcal{X}}^{2}$.
To solve these types of problems, one popular approach includes penalty based methods which are described as differential inclusion
$\dot{x}(t)+A(x(t))+\beta(t)\partial f(x(t))\ni 0$ (24)
where the penalty parameter $\beta(t)$ tends to infinity. As demonstrated in [9], when the monotone operator is a sub-differential $A=\partial g$, then we may equivalently consider the differential inclusion
$\dot{x}(t)+\lambda(t)\partial g(x(t))+\partial f(x(t))\ni 0$
with vanishing parameter $\lambda(t)$. In summary, analyses of the above differential inclusion can be translated to the differential equation
$\dot{x}(t)+\nabla f(x(t))+\lambda(t)x(t)=0$ (25)
describing the regularized steepest descent in continuous time. Note that reg-GD defined in (3) can be interpreted as explicit Euler discretization of (25).
D.2.1 Related work in the deterministic setting
The analysis of dynamical systems corresponding to (24) with $\partial f=0$ dates back to the 1970s. For instance, in [13], it was shown that for $A=\partial g$, where $g$ is lower semicontinuous, proper, and convex, the trajectory converges weakly to a minimizer of $g$. More generally, for maximal monotone operators $A$, the ergodic average of the trajectory converges weakly to a point in $A^{-1}(\{0\})$ [23].
The penalty-based differential inclusion (24) was introduced in [9], where the authors established weak ergodic convergence (and even strong convergence for strongly monotone operators $A$) under the integrability condition
$\int_{0}^{\infty}\beta(t)\Big[\Psi^{\ast}\big(\frac{p}{\beta(t)}\big)-\sigma_{C}\big(\frac{p}{\beta(t)}\big)\Big]\,\mathrm{d}t<\infty\quad\text{for all}\ p\in\mathrm{range}(N_{C})\,$,
where $\Psi^{\ast}$ denotes the Fenchel conjugate of $\Psi$, and $\sigma_{C}$ is the support function of the set $C$. This condition is now commonly referred to as the Attouch–Czarnecki condition.
Note that a similar condition arises in our analysis as the final requirement in (7). While our condition can be characterized via the Łojasiewicz inequality, the Attouch–Czarnecki condition can be characterized using a quadratic error bound of the form
$\Psi(x)\geq C\,\mathrm{dist}(x,C)^{2}\,$,
which implies that
$\Psi^{\ast}(p)-\sigma_{C}(p)\leq\frac{\|p\|^{2}}{2C}\,$,
see for instance [9, 11] for more details. In this case, the Attouch–Czarnecki condition is guaranteed under integrability conditions on the penalty function $\beta(\cdot)$.
In the discrete-time setting, the Attouch–Czarnecki condition translates into a summability condition involving both the penalty sequence and the step-sizes. For instance, [60] introduces a coupled gradient method with exterior penalization, leading to the condition
$\sum_{n\in\mathbb{N}}\alpha_{n}\beta_{n}\Big[\Psi^{\ast}\big(\frac{p}{\beta_{n}}\big)-\sigma_{C}\big(\frac{p}{\beta_{n}}\big)\Big]<\infty\,.$
Here, the author considers the case where $A=\nabla g$ and both $\nabla f$ and $\nabla g$ are Lipschitz continuous, establishing weak convergence under convexity of $g$, and strong convergence when $g$ is strongly convex. Other results in the discrete-time setting include splitting-based discretization schemes [10, 11, 27] whose convergence analysis rely on some similar variant of the Attouch–Czarnecki condition.
D.2.2 Convergence rate for reg-GD
@@
In the following, we quantify the rate of convergence of reg-GD defined in (3). Our derived rates are consistent with known results for the ODE (25). In particular, for sufficiently small step-sizes, i.e. $\alpha_{k}\leq\frac{2}{L}$, our results match those derived in [7, Theorem 5] when defining the numerical time $t_{k}=\sum_{i=1}^{k}\alpha_{i}$ for $k\in\mathbb{N}$ and noting that $t_{k}\sim\frac{C_{\alpha}}{1-q}k^{1-q}$ and $\lambda_{k}\sim C_{\lambda}(\frac{1-q}{C_{\alpha}}t_{k})^{-p/(1-q)}$ for $q<1$. The proof follows the strategy of [7, Theorem 5].
Theorem D.5.
@@
Suppose that Section˜1.1 is satisfied. Let $C_{\alpha},C_{\lambda}>0$, $p\in(0,1]$ and $q\in[0,1-p]$. Let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (3) for all $k\in\mathbb{N}$, $(\lambda_{k})_{k\in\mathbb{N}}=(C_{\lambda}k^{-p})_{k\in\mathbb{N}}$ and $(\alpha_{k})_{k\in\mathbb{N}}=(C_{\alpha}k^{-q})_{k\in\mathbb{N}}$ such that the following conditions are satisfied:
@@
(i): Appendix˜D with $D_{k}\equiv 0$ guarantees that, for all $k\in\mathbb{N}_{0}$ with $\alpha_{k}\leq\frac{2}{L+\lambda_{k}}$, one has
@@
Set $\beta=1-q$ and for $k\in\mathbb{N}$ define [^1] In order to avoid confusion, we note that $\varphi_{k}$ is defined differently as in the proofs of Theorem D.3 and Theorem D.4 $\varphi_{k}=E_{k}k^{\beta}$. By assumption on $(\alpha_{k})_{k\in\mathbb{N}}$ one has $\alpha_{k}<\frac{2}{L}$ for all but finitely many indices $k$. Therefore there exists an $N\in\mathbb{N}$ such that for all $k>N$
@@
where in the case q=1pq=1-p and $q\neq 0$ we have used that $2C_{\lambda}C_{\alpha}>1-q$ and in the case $q=0$ and $p=1$ we have used that $2C_{\lambda}C_{\alpha}(1-\mbox{$\frac{LC_{\alpha}}{2}$})>1$. Inserting these inequalities in (26),
@@
(ii): Follows from (i) and Appendix˜D, using that $p\leq 1-q$.
(iii): Follows from (i) and Appendix˜D.
@@
In the spirit of Section˜2.3, we will derive optimal decay rates for the step-size and regularization decay for the convergence to the minimum-norm solution under the additional assumption that there exist $C_{\text{reg}},\xi>0$ with
@@
see also Section˜3. Using Theorem˜D.5, one has
@@
In Figure˜14, we illustrate the convergence rate on depending on the decay-rates $p,q$ for $\xi=\frac{1}{4}$.
Figure 14: Convergence rate for $\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}$ in the situation of Theorem˜D.5 under the Polyak-Łojasiewicz inequality, i.e. under (10) with $\xi=\frac{1}{4}$.
D.3 $L^{2}$-convergence rate for reg-SGD: Proof of Theorem˜2.3
In the following, we formulate Theorem˜2.3 in more details and provide a full prove.
Theorem D.6.
Suppose that Section˜1.1 and Section˜1.1 are satisfied. Let $C_{\alpha},C_{\lambda}>0$, $p\in(0,\frac{1}{2}]$ and $q\in(p,1-p]$. Let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with $(\alpha_{k})_{k\in\mathbb{N}}=(C_{\alpha}k^{-q})_{k\in\mathbb{N}}$ and $(\lambda_{k})_{k\in\mathbb{N}}=(C_{\lambda}k^{-p})_{k\in\mathbb{N}}$. If q=1pq=1-p we additionally assume that $2C_{\lambda}C_{\alpha}>1-q$. Then, it holds that $\lim_{k\to\infty}\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]=0$ and
@@
We will only prove the first claim (i). Due to Theorem˜2.2, one has $\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]\to 0$ if $q>p$ and $p+q<1$. The statements (ii)-(iii) follow analogously to the proof of Theorem˜D.5. Let $\beta=\min(1-q,q-p)$ and $\varphi_{k}=\mathbb{E}[E_{k}]k^{\beta}$. By assumption on $(\alpha_{k})_{k\in\mathbb{N}}$ one has $\alpha_{k}<\frac{2}{L}$ for all but finitely many $k\in\mathbb{N}$ so that, using Appendix˜D,
@@
Since $q>p$, $p+q\geq 1$, and $2C_{\lambda}C_{\alpha}>\min(p,1-2p)$ in the case that q=1pq=1-p one can show as in (27) that there exist $\varepsilon^{\prime}>0$ and $N\in\mathbb{N}$ such that for all $k\geq N$
@@
By choice of $\beta$, one has $p+q=\max(1+p-\beta,2q-\beta)$. Therefore, we can show analogously to the proof of Theorem˜D.5 that
@@
D.4 Almost sure convergence rate for reg-SGD: Proof of Theorem˜2.4
In the following, we formulate Theorem˜2.4 in more details and provide a full prove. The proof of the almost sure convergence rates requires a sophisticated application of the Robbins-Siegmund theorem, Appendix˜B. For this, we use the variation of constants formula to separate the influence of the stochastic noise term $(D_{k})_{k\in\mathbb{N}}$ and the deterministic change in the global minimum of the regularized objective function $(x_{\lambda_{k}}-x_{\lambda_{k+1}})_{k\in\mathbb{N}}$.
Theorem D.7.
Suppose that Section˜1.1 and Section˜1.1 are satisfied. Let $C_{\alpha},C_{\lambda}>0$, $p\in(0,\frac{1}{2})$ and $q\in(\frac{1}{2},1-p]$. Let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with $(\alpha_{k})_{k\in\mathbb{N}}=(C_{\alpha}k^{-q})_{k\in\mathbb{N}}$ and $(\lambda_{k})_{k\in\mathbb{N}}=(C_{\lambda}k^{-p})_{k\in\mathbb{N}}$. Let $\beta\in(0,2q-1)$ and, if q=1pq=1-p, we assume that $2C_{\lambda}C_{\alpha}>\min(\beta,1-q)$. Then,
@@
We will only prove property (i). Properties (ii)-(iv) follow analogously to the proof of Theorem˜D.5. By Appendix˜D, for all $k\in\mathbb{N}$ with $\alpha_{k}\leq\frac{2}{L+\lambda_{k}}$ one has
@@
Let $\beta\in(0,2q-1)$, $\tilde{\beta}=\min(\beta,1-q)$ and $\varphi_{k}=\tilde{E}_{k}k^{\tilde{\beta}}$. By assumption on $(\alpha_{k})_{k\in\mathbb{N}}$ one has $\alpha_{k}<\frac{2}{L}$ for all but finitely many $k\in\mathbb{N}$ so that
$\mathbb{E}[\varphi_{k}\mid\mathcal{F}_{k-1}]$ $\leq\Big(1-2\lambda_{k}\alpha_{k}\Big(1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big)+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}A\Big)\frac{k^{\tilde{\beta}}}{(k-1)^{\tilde{\beta}}}\varphi_{k}$
@@
Since $q>p$, $p+q\geq 1$, and $2C_{\lambda}C_{\alpha}>\tilde{\beta}$ in the case that q=1pq=1-p, one can show as in (27) that there exist $\varepsilon^{\prime}>0$ and $N\in\mathbb{N}$ such that for all $k\geq N$
@@
[…]
$\Psi_{k}=\Big(1-2\lambda_{k}\alpha_{k}\Big(1-\frac{L+\lambda_{k}}{2}\alpha_{k}\Big)+\frac{L+\lambda_{k}}{2}\alpha_{k}^{2}A\Big)\Psi_{k-1}+\frac{\lambda_{k}-\lambda_{k+1}}{2}\|x_{\ast}\|_{\mathcal{X}}^{2}.$
[…]
Appendix E Properties of the Tikhonov regularization